OpenStudy (anonymous):

Solve x – 2y = 6 3x – 6y = 18

OpenStudy (anonymous):

Y = (x-6)/2

OpenStudy (anonymous):

x = 2y+6 3(2y+6) - 6y = 18 6y + 18 - 6y = 18 the 6ys cancel out 18 = 18 all points on line are solution, its the same line